2. value: 10.00 points Milano Pizza is a small neighborhood pizzeria that has a small area...

2.

value:
10.00 points

Milano Pizza is a small neighborhood pizzeria that has a small area for in-store dining as well as offering take-out and free home delivery services. The pizzeria’s owner has determined that the shop has two major cost drivers—the number of pizzas sold and the number of deliveries made. Data concerning the pizzeria’s costs appear below:

  

Fixed Cost
per Month
Cost per
Pizza
Cost per
Delivery
  Pizza ingredients $ 4.70   
  Kitchen staff $ 5,970   
  Utilities $ 640    $ 0.60   
  Delivery person $ 3.40   
  Delivery vehicle $ 660    $ 1.80   
  Equipment depreciation $ 424   
  Rent $ 1,930   
  Miscellaneous $ 760    $ 0.06   

  

       In November, the pizzeria budgeted for 1,650 pizzas at an average selling price of $18 per pizza and for 250 deliveries.
       Data concerning the pizzeria’s operations in November appear below:

  

Actual
Results
  Pizzas 1,750  
  Deliveries 230  
  Revenue $ 32,080  
  Pizza ingredients $ 7,750  
  Kitchen staff $ 5,910  
  Utilities $ 900  
  Delivery person $ 782  
  Delivery vehicle $ 992  
  Equipment depreciation $ 424  
  Rent $ 1,930  
  Miscellaneous $ 808  


Required:
1.

Complete the flexible budget performance report that shows both revenue and spending variances and activity variances for the pizzeria for November. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

     

rev: 08_14_2014_QC_51991
Answer #1
Static Budget
Income
Sales Revenue $            29,700 =1650*18
Expenses
Pizza Ingredients $               7,755 =1650*4.7
Kitchen Staff $               5,970 5970
Utilities $               1,630 =1650*0.6+640
Delivery Person $                  850 =250*3.4
Delivery Vehicle $               1,110 =250*1.8+660
Equipment Depreciation $                  424 424
Rent $               1,930 1930
Misc Costs $                  859 =1650*0.06+760
Total Expenses $            20,528
Net Income $               9,172
Flexible Budget
Income
Sales Revenue $            31,500 =1750*18
Expenses
Pizza Ingredients $               8,225 =1750*4.7
Kitchen Staff $               5,970 5970
Utilities $               1,690 =1750*0.6+640
Delivery Person $                  782 =230*3.4
Delivery Vehicle $               1,074 =230*1.8+660
Equipment Depreciation $                  424 424
Rent $               1,930 1930
Misc Costs $                  865 =1750*0.06+760
Total Expenses $            20,960
Net Income $            10,540
Actual Results Flexible Budget Activity Variance
Income
Sales Revenue $            32,080 $     31,500 $               580 F
Expenses
Pizza Ingredients $               7,750 $       8,225 $             -475 F
Kitchen Staff $               5,910 $       5,970 $               -60 F
Utilities $                  900 $       1,690 $             -790 F
Delivery Person $                  782 $           782 $                  -  
Delivery Vehicle $                  992 $       1,074 $               -82 F
Equipment Depreciation $                  424 $           424 $                  -  
Rent $               1,930 $       1,930 $                  -  
Misc Costs $                  808 $           865 $               -57 F
Total Expenses $            19,496 $     20,960 $         -1,464 F
Net Income $            12,584 $     10,540 $           2,044 F
Actual Results Static Budget Spending Variance
Income
Sales Revenue $            32,080 $     29,700 $           2,380 F
Expenses
Pizza Ingredients $               7,750 $       7,755 $                  -5 F
Kitchen Staff $               5,910 $       5,970 $               -60 F
Utilities $                  900 $       1,630 $             -730 F
Delivery Person $                  782 $           850 $               -68 F
Delivery Vehicle $                  992 $       1,110 $             -118 F
Equipment Depreciation $                  424 $           424 $                  -  
Rent $               1,930 $       1,930 $                  -  
Misc Costs $                  808 $           859 $               -51 F
Total Expenses $            19,496 $     20,528 $         -1,032 F
Net Income $            12,584 $       9,172 $           3,412 F